OpenStudy (anonymous):

find the inverse function of f(x)=e^( 10(x-10) )

jimthompson5910 (jim_thompson5910):

Replace f(x) with y, and then swap x and y to get x = e^(10(y-10)) Now solve for y x = e^(10(y-10)) ln(x) = 10(y-10) ln(x) = 10y-100 ln(x)+100 = 10y (1/10)(ln(x)+100) = y y = (1/10)(ln(x)+100) y = ln(x)/10+10 So the inverse function is g(x) = ln(x)/10+10

OpenStudy (anonymous):

is this \[e^{10x-100}\]

OpenStudy (anonymous):

yes that's the question

OpenStudy (anonymous):

then jim thompson has it

OpenStudy (anonymous):

ummm, haven't learned the "In" thing yet..what's that?

OpenStudy (anonymous):

inverse of \[e^x\]

jimthompson5910 (jim_thompson5910):

the ln stands for the natural log. It's just a logarithm with base 'e'.

jimthompson5910 (jim_thompson5910):

In other words, \[ln(x)=\log_{e}(x)\]

OpenStudy (anonymous):

ooo icic thank you so much

OpenStudy (anonymous):

if you haven't learned about log, this problem will make no sense

OpenStudy (anonymous):

can u write in the log form to show me?

jimthompson5910 (jim_thompson5910):

Inverse function is \[f^{-1}(x)=\frac{\log_{e}(x)}{10}+10\] using common log notation.

OpenStudy (anonymous):

ok this is how i did x = 10(y-10)log e x = 10ylog e- 100log e x = 10( y log e- 10log e) x/10 = y log e- 10log e (x/10) + 10log e= ylog e x/(10log e) +10=y

OpenStudy (anonymous):

can u check what I did wrong there? i just couldn't get the same answer like urs

jimthompson5910 (jim_thompson5910):

When you took the log of the right side, you forgot to take the log of the left side as well. Also, the log e on the right side will turn into 1, which means that it essentially goes away.